A triple-X (XXX) individual has how many Barr bodies? O 2 O 3 04
Answer #1

The number of Barr bodies present in a xxx(triple x) individual is : 2

Number of Bam bodies = (3-1) whene na neember of x chromosome for xxx indivisual n=B So, no. of barors bodies = (n-1)= 3-1=2​​​​

Barr bodies are the theme activated X chromosomes formed in somatic cells during embryonic development. Once the inactivation of course for the all the cells will have the same X chromosome in activated.

In this case there are three X chromosomes out of which 2 X chromosomes are inactivated and converted into Barr bodies.